if xy=4 prove that x(dy/dx+y^2)=3y

Grade:12

2 Answers

bharat bajaj IIT Delhi
9 years ago
Xy = 4
Differntiate wrt x
x.dy/dx + y = 0
dy/dx = -y/x

LHS
x(dy/dx + y^2)
= x(-y/x + y^2)
= x(-4/x^2 + 16/x^2)
= 12/x = 3. 4/x
= 3y = RHS

Thanks
Bharat Bajaj
Askiitians faculty

Arun Kumar IIT Delhi
9 years ago
Hi

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty